# SF-Loc: A Visual Mapping and Geo-Localization System based on Sparse Visual Structure Frames

### v2

@@ Abstract
For high-level geo-spatial applications and intelligent robotics, accurate global pose information is of crucial importance. Map-aided localization is a universal approach to overcome the limitations of global navigation satellite system (GNSS) in challenging environments. However, current solutions face challenges in terms of mapping flexibility, storage burden and re-localization performance. In this work, we present SF-Loc, a lightweight visual mapping and map-aided localization system, whose core idea is the map representation based on sparse frames with dense but compact depth, termed as visual structure frames. In the mapping phase, multi-sensor dense bundle adjustment (MS-DBA) is applied to construct geo-referenced visual structure frames. The local co-visbility is checked to keep the map sparsity and achieve incremental mapping. In the localization phase, coarse-to-fine vision-based localization is performed, in which multi-frame information and the map distribution are fully integrated. To be specific, the concept of spatially smoothed similarity (SSS) is proposed to overcome the place ambiguity, and pairwise frame matching is applied for efficient and robust pose estimation. Experimental results on the cross-season dataset verify the effectiveness of the system. In complex urban road scenarios, the map size is down to 3 MB per kilometer and stable decimeter-level re-localization can be achieved.
Index Terms:
Dense bundle adjustment, multi-sensor fusion, visual localization and mapping, place recognition
@@ I Introduction
Global localization is a critical problem in robotics and autonomous driving applications [1, 2]. For outdoor scenarios, global navigation satellite system (GNSS) is a commonly used technology for achieving global drift-free localization[3]. However, precise GNSS generally relies on high-precision services and products, while low-cost GNSS can hardly provide stable and accurate localization solutions in complex environments[4]. Map matching is another method for achieving global localization. By associating observations of the surrounding environment from external sensors widely equipped on intelligent devices (such as cameras) with a priori maps, different levels of global localization can be achieved, depending on the map representation[5, 6].
To achieve the practicality of map-aided localization, both the mapping and the re-localization phases should be carefully addressed. As to the mapping phase, the efficiency of map construction, the accuracy of the map elements and the storage burden of map data are some of the key factors. Considering the need of incremental mapping and crowd-sourced map updating, the flexibility of the map representation is also crucial. Current map forms, like vector map[7], 3D point cloud[8, 9], occupancy grid[10] and neural representations[11, 12], still cannot meet the above requirements perfectly, especially when considering vision-based schemes.
As to re-localization, the system functionality is highly coupled with the map representation. High recall and high localization accuracy generally need more detailed and distinguishable information of the environment, which seems contradictory with the expectation of a lightweight map. How to choose a practical map representation that meets the need of efficient mapping and effective re-localization, and develop a pipeline properly adapted to the representation, is the main concern of this work.
@@
In this work, we propose SF-Loc, a lightweight visual mapping and map-aided localization system based on sparse visual structure frames. The visual structure frame is defined as a geo-tagged image frame with compressed RGB map and full-view depth map, which itself supports standalone and flexible re-localization. The word “structure” comes from two aspects, as the frame stores the environmental structure information (in the depth form), and the map frames are sparsely structured with special consideration for efficient association and reducing redundancy. Based on the concept, SF-Loc provides a pipeline for the auto-generation of visual structure frames through multi-sensor dense bundle adjustment (MS-DBA), as well as an implementation for coarse-to-fine map-based localization. The contributions of this work are listed as follows:
1) We provide a pipeline for large-scale mapping based on the visual-structure-frame map representation, which is built upon the integration of deep dense bundle adjustment and multi-sensor factor graph optimization (FGO).
2) We provide a coarse-to-fine map-aided localization pipeline which fully utilizes deep place recognition, local feature matching and spatial-temporal information association.
3) Real-world, cross-season experiments are conducted to validate different phases of the system.
4) The code is made open-source111 https://github.com/GREAT-WHU/SF-Loc to benefit the community.
@@ II-A VSLAM with multi-sensor fusion
Visual SLAM, which has been extensively studied in the past decade, is considered an effective and low-cost solution for relative pose estimation and mapping. Considering the form of visual information, point features[8, 13], line features[14] and direct photometric errors[15] have been exploited. By integrating IMU[16, 17, 18], GNSS[19, 3], and other sensors[20], the practicality of VSLAM can be improved in terms of robustness, continuity, accuracy and geo-referencing functionality. On this basis, multi-session and multi-agent SLAM systems are developed[21, 22]. Learning-based VSLAM frameworks are also investigated in recent implementations, covering neural representations[23, 24], data-driven association[25, 26, 27] and end-to-end pipelines[28, 29].
For map reuse of multi-sensor integrated VSLAM, obtaining globally smoothed pose estimation in the geographic coordinate system is important, as it impacts the absolute accuracy of the map and the practicality of user-side localization.
@@ II-B Visual place recognition
Visual place recognition (VPR) is a long-existing problem, which plays a crucial role in loop closure, multi-session mapping, re-localization and so on. Early approaches use handcrafted models (like bag-of-words[30] and VLAD[31]) to aggregate local feature descriptors, thus to evaluate the scene similarity. To overcome the long-term change challenge, sequential information is utilized[32].
In the last decade, deep neural network (DNN)-based methods, like NetVLAD[33], have greatly improved the practicality and accuracy of VPR. Following this trend, TransVPR[34] uses the transformer to incorporate multi-level attention mechanisms and improve the VPR performance. CosPlaces[35] transforms the place recognition problem from contrastive learning to a classification problem, which enables efficient training on 10-million-level large-scale datasets and is further enhanced with better viewpoint robustness[36]. AnyLoc[37] utilizes the visual foundation model to achieve ubiquitous VPR for different scenarios. It is noted that, most existing methods don’t consider comprehensively the spatial distribution of database images and user-side relative poses, which could serve as powerful information to improve the stability of VPR.
@@ II-C Fine-grained visual re-localization
For high-accuracy visual re-localization (e.g. decimeter- or centimeter-level), it is generally needed to perform feature- or instance-level association between user perception and the map elements. Many of these methods follow a coarse-to-fine pipeline. In ORB-SLAM3[8] and VINS-Mono[16], re-localization is performed through bag-of-words-based place recognition and pointwise matching between new feature observations and the point cloud map. Such pipeline is further improved in HLoc[9], where the power of deep learning is fully utilized in both mapping and localization phases. Besides, monocular depth prediction is exploited in [39] to achieve map-free visual re-localization.
Along with these methods, other approaches try to exploit higher-level information for map matching. Instance-level object modeling is integrated into classic VSLAM systems to overcome viewpoint variation and achieve long-term consistency[40, 41, 42, 43]. In recent works, bird’s-eye view (BEV) features are utilized to associate with priori vector maps to achieve accurate pose estimation, both for outdoor[44] and indoor scenarios[45]. The tradeoff of semantic-based re-localization is the loss of low-level, sometimes subtle but distinguishable, visual cues, which might lead to higher ambiguity.
@@ III System Overview
The proposed system is built upon the concept of the visual-structure-frame-based map representation, and is mainly divided into the mapping phase and the localization phase, as shown in Fig. 2.
In the mapping phase, multi-sensor information is used to generate the feature-free map frames with recovered dense depth. Global co-visibility graph optimization is performed periodically in a low frequency to achieve global optimal pose estimation. On this basis, the local co-visibility is checked to keep the map sparsity and to achieve incremental mapping. Besides, a DNN-based VPR model is employed to compute the global descriptors of map frames.
In the localization phase, deep visual place recognition and feature matching models are exploited to achieve coarse-to-fine localization. During this process, multi-frame information and the map frame distribution are fully considered to improve the recall rate and the accuracy.
One principle behind the proposed map representation is to keep the visual information raw and non-encoded. The first advantage of doing so is that it ensures users have flexibility when using the map, without being constrained by any intermediate forms of information. The second advantage is that, the compression of a raw image is effective by applying mature techniques like JPEG[46], without losing the dominant geometric or appearance information.
@@ IV Mapping Phase
In this phase, the implementation details of the mapping phase are presented.
@@ IV-A Multi-Sensor Dense Bundle Adjustment
The concept of dense bundle adjustment (DBA) is proposed in[26], which is inherited in this work. Specifically, sequential images are fed into a convolutional neural network (CNN) with gated recurrent units (GRUs) to compute dense optical flows among co-visible image pairs. To explain this, consider the rigid flow produced by dense projection of one image pair (source frame $i$ projected to target frame $j$)
@@
where $\Pi_{c}$ is the camera projection model, $\mathbf{T}_{ij}$ is the relative camera pose, $\mathbf{u}_{i}$ denotes the $n=(H/8)\times(W/8)$ grid-like 2D points in the image frame, $\bm{\lambda}_{i}$ is the inverse depth map.
The optical flow module takes the CNN-encoded image feature map and the rigid flow as input and outputs the corrected optical flow (in the residual form) together with the weight $\mathbf{w}_{ij}$, which leads to the following error equation
[…]
where $\bm{\xi}_{i}$, $\bm{\xi}_{j}$ are Lie algebras of the camera poses, $\mathbf{J}_{i}$, $\mathbf{J}_{j}$, $\mathbf{J}_{\bm{\lambda}_{i}}$ are Jacobians derived from (1).
The residual-form dense optical flows are treated as re-projection errors, which are further used for bundle adjustment to simultaneously estimate the poses and dense depths of the images. Furthermore, the estimated poses and depths are recurrently fed back into the optical flow module, thus to iteratively refine all the geometric information. This process is end-to-end trainable, which shows good generalization in real-world scenes with synthetic training[26].
In our early work[48], we demonstrate the possibility of tightly integrating DBA with multi-sensor information. To be specific, the Hessian form of (2) is derived and stacked for all projections from frame $i$, leading to
$\left[\begin{matrix}\mathbf{v}_{i}\\ \mathbf{z}_{i}\end{matrix}\right]=\left[\begin{matrix}\mathbf{B}_{i}&\mathbf{E}_{i}\\ \mathbf{E}^{\top}_{i}&\mathbf{C}_{i}\end{matrix}\right]\left[\begin{matrix}\bm{\xi}_{i,1,2,\dots,N}\\ {\delta\bm{\lambda}_{i}}\end{matrix}\right]$ (3)
where $\bm{\xi}_{i,1,\cdots,N}$ is the stacked Lie algebras of camera poses.
@@
$(\underbrace{\mathbf{B}_{i}-\mathbf{E}_{i}\mathbf{C}_{i}^{-1}\mathbf{E}_{i}^{\top}}_{\mathbf{H}_{c,i}})\bm{\xi}_{i,1,\cdots,N}=\underbrace{\mathbf{v}_{i}-\mathbf{E}_{i}\mathbf{C}_{i}^{-1}\mathbf{z}_{i}}_{\mathbf{v}_{c,i}}$ (4)
in which the inversion operation of $\mathbf{C}_{i}$ is efficient as it is a diagonal matrix.
@@
The Hessian-form pose constraint, which embeds the DBA information of co-visible images, is fed to a sliding-window factor graph based on GTSAM[47] for multi-sensor fusion, as shown in Fig. 3. A generic graph design is employed, which allows for flexible integration of different kinds of sensors or information. The states in the graph are as follows
$\mathbf{x}_{k}=\left(\begin{matrix}\mathbf{T}^{w}_{b_{k}}&\mathbf{v}^{w}_{b_{k}}&\mathbf{b}_{k}\end{matrix}\right),k\in\mathcal{K}$ (5)
$\mathbf{T}^{w}_{b_{k}}=\left[\begin{matrix}\mathbf{R}^{w}_{b_{k}}&\mathbf{t}^{w}_{b_{k}}\\ 0&1\end{matrix}\right]\in SE(3),\ \ \mathbf{b}_{k}=\left(\begin{matrix}\mathbf{b}_{a,k}&\mathbf{b}_{g,k}\end{matrix}\right)$ (6)
where $\mathcal{K}$ is the frame set of the sliding window, $\mathbf{T}^{w}_{b_{k}}$ is the body (IMU) poses in the world frame, $\mathbf{v}^{w}_{b_{k}}$ is the velocity, $\mathbf{b}_{a,k}$ and $\mathbf{b}_{g,k}$ are accelerometer/gyroscope biases.
Without losing generality, the following factors are considered in this work:
@@
$\mathbf{E}_{c}(\mathbf{x}_{c})=\frac{1}{2}\mathbf{x}_{c}^{\top}\mathbf{H}_{c}\,\mathbf{x}_{c}-\mathbf{x}_{c}^{\top}\mathbf{v}_{c}$ (7)
where $\mathbf{H}_{c}$, $\mathbf{v}_{c}$ are the Hessian matrix and vector of DBA (stacked from $\mathbf{H}_{c,i}$, $\mathbf{v}_{c,i}$) with all the depth states eliminated, $\mathbf{x}_{c}$ is the related camera poses. To associate the camera poses with the IMU-centered pose states, the IMU-camera extriniscs are used to transform (7), as detailed in[48].
@@
where $\tilde{\mathbf{p}}^{n}_{g}$ is the measured position of the GNSS phase center in the navigation frame, $\mathbf{t}^{b}_{g}$ is the IMU-GNSS lever-arm, $\mathbf{T}^{n}_{w}$ is a fixed world-to-navigation transformation obtained through the initial alignment.
Probabilistic marginalization is applicable to the above factor graph, following the common operation based on Schur complement[17]. For DBA information, a sub-graph related to the oldest frame is temporarily constructed to generate the marginalization information. Considering the marginalization term, the overall cost function of the sliding-window factor graph is defined as follows
$\begin{split}\min\ \ \mathbf{E}_{m}(\mathbf{x})+\mathbf{E}_{c}(\mathbf{x}_{c})&\\ +\sum_{k\in\mathcal{K}}\left\|\mathbf{r}_{b}(\mathbf{x}_{k},\mathbf{x}_{k+1})\right\|^{2}_{\bm{\Omega}_{b}}&+\sum_{k\in\mathcal{G}}\rho_{C}\left(\left\|\mathbf{r}_{g}(\mathbf{x}_{k})\right\|^{2}_{\bm{\Omega}_{g}}\right)\end{split}$ (10)
where $\mathbf{E}_{m}(\mathbf{x})$ is the marginalization term in the quadric cost function form, $\bm{\Omega}_{b}$ and $\bm{\Omega}_{g}$ are uncertainties of the factors, $\rho_{C}(\cdot)$ is the Cauchy loss function.
@@
$\delta\bm{\lambda}_{i}=\mathbf{C}_{i}^{-1}(\mathbf{z}_{i}-\mathbf{E}_{i}^{\top}\bm{\xi}_{i,1,\dots,N})$ (11)
then the estimated poses and depths are fed back into the optical flow module to refine the optical flows and adjust the weights. For every projection, the optical flow module will be recurrently called for several times until mature. The overall process simultaneously adjust the pose estimation, the dense data association and the depths, trying to achieve consistency of the above information.
An issue that should be considered is the obtainment of global optimal pose estimation. This is important both for monocular visual-inertial odometry (due to the dynamic-dependent observability) and GNSS-integrated navigation (due to intermittent GNSS availability), which can’t be simply achieved in a sliding-window estimator. In our system, we maintain a global factor graph, which collects the marginalized factors from the sliding-window factor graph. The objective of the global factor graph is written as follows
@@
where $\mathbf{E}_{c,k}(\mathbf{x}_{c,k})$ is the DBA information that takes frame $k$ as the source frame, $\mathbf{x}_{c,k}$ is the related camera poses. Each $\mathbf{E}_{c,k}$ reflects the local co-visibility around frame $k$.
@@
Figure 4: Illustration of the co-visibility checking. (a) Co-visibility for one image pair; (b) Local co-visibility checking, the dashed-line frame won’t be inserted due to significant co-visibility with one nearby frame; (c) An example of the sparsified structure frame map.

### v3

@@ Abstract
For high-level geo-spatial applications and intelligent robotics, accurate global pose information is of crucial importance. Map-aided localization is a generally applicable approach to overcome the limitations of global navigation satellite system (GNSS) in challenging environments. However, current solutions face challenges in terms of mapping efficiency, storage burden and re-localization performance. In this work, we present SF-Loc, a lightweight visual mapping and map-aided localization system, which is built upon the image-based map representation with dense but compact depth, termed as visual structure frames. In the mapping phase, multi-sensor dense bundle adjustment (MS-DBA) with global optimization is applied to construct geo-referenced visual structure frames. Besides, a co-visibility-based mechanism is developed to achieve incremental mapping and keep the map sparsity. In the localization phase, coarse-to-fine vision-based localization is performed, in which multi-frame information and the map distribution are fully leveraged. To be specific, spatiotemporally associated similarity (SAS) is proposed to overcome the place ambiguity, and pairwise frame matching is applied for robust pose estimation. Experimental results on the real-world cross-temporal dataset verify the effectiveness of the system. In complex urban road scenarios, the map size is down to 3 MB per kilometer and stable decimeter-level re-localization can be achieved.
@@ I Introduction
Global localization is a critical problem in robotics and autonomous driving applications [1, 2]. For outdoor scenarios, global navigation satellite system (GNSS) is a commonly used technology for achieving global drift-free localization[3]. However, precise GNSS generally relies on high-precision services and products, while low-cost GNSS can hardly provide stable and accurate localization solutions in complex environments[4]. Map matching is another method for global localization. By associating observations from exteroceptive sensors (such as cameras) mounted on intelligent devices with the prior maps, different levels of global localization can be achieved, depending on the map representation[5, 6].
To achieve the practicality of map-aided localization, both mapping and re-localization need to be handled. As to the mapping phase, the efficiency of map construction, the map accuracy, the storage burden and the information sufficiency for localization are some of the key factors. Current map forms, like vector map[7], 3D point cloud[8, 9], occupancy grid[10] and neural representations[11, 12], still cannot fully satisfy all these requirements simultaneously, especially when considering vision-based low-cost schemes.
As to re-localization, the system functionality is highly coupled with the map data. High recall and high localization accuracy generally need more detailed and distinguishable information of the environment, which seems contradictory with the expectation of a lightweight map.
It is noted that, the image itself can serve as an advantageous landmark for localization-oriented maps, which is rich in cues for fine-grained association, widely accessible on ubiquitous devices, and free from cross-modality information loss. However, the following issues should be considered:
  1. 1.
    How to efficiently and accurately recover the absolute poses of images at a large scale?
  2. 2.
    Since images are storage-intensive, to what extent can such image maps be compressed (or sparsified) while still meeting user localization requirements?
  3. 3.
    What auxiliary information is needed for accurate localization? And how to achieve high-recall localization in cross-temporal or weak-recognizability scenarios?
@@
In this work, we propose SF-Loc (as shown in Fig. 1), a lightweight visual mapping and map-aided localization system based on sparse visual structure frames, aiming to comprehensively address the above issues. The visual structure frame is defined as a geo-tagged image frame with full-view depth, which itself supports standalone and flexible re-localization. Based on the concept, SF-Loc provides a holistically designed framework for efficient, accurate mapping and high-availability re-localization. The contributions of the paper are as follows:
1) We provide a pipeline for large-scale mapping based on the visual-structure-frame map representation, which extracts accurate geometric information from global optimized multi-sensor dense bundle adjustment (MS-DBA) and supports the generation (incremental, multi-session, etc.) of practical localization-oriented maps.
2) We provide a coarse-to-fine map-aided localization pipeline based on the structure frame map, which leverages the spatiotemporally associated similarity (SAS) and pairwise frame matching for high-recall localization.
3) Real-world, cross-temporal experiments are conducted to validate different phases of the system.
4) The code is made open-source [^1] https://github.com/GREAT-WHU/SF-Loc to benefit the community.
@@ II-A VSLAM with multi-sensor fusion
Visual simultaneous localization and mapping (SLAM), which has been extensively studied in the past decade, is considered an effective and low-cost solution for relative pose estimation and mapping. By integrating IMU[13, 14], GNSS[3], and other sensors[15], the practicality of VSLAM can be improved in terms of robustness, continuity, accuracy and geo-referencing functionality. On this basis, multi-session and multi-agent SLAM systems are developed[16, 17]. Learning-based VSLAM frameworks are also investigated in recent implementations, covering neural/learnable representations including neural radiance fields (NeRF)[18] and 3D Gaussian splatting (3DGS)[19], data-driven association[20, 21, 22] and end-to-end pipelines[23]. However, to extract a practical reusable map from multi-sensor integrated VSLAM is still an open problem, especially when considering absolute accuracy and cross-temporal functionality.
@@ II-B Visual place recognition
Visual place recognition (VPR) is a long-existing problem, which plays a crucial role in loop closure, multi-session mapping, re-localization and so on. Early approaches use handcrafted models (like bag-of-words[24] and VLAD[25]) to aggregate local feature descriptors, thus to evaluate the place similarity. To overcome the long-term change challenge, sequential information is utilized[26].
In the last decade, deep neural network (DNN)-based methods, like NetVLAD[27] and its successors with stronger backbones[28, 45], have greatly improved the practicality and accuracy of VPR. CosPlaces[29] models place recognition as classification, which enables large-scale training and is further enhanced with better viewpoint robustness[30]. AnyLoc[31] utilizes the visual foundation model to achieve ubiquitous VPR for different scenarios. It is noted that, most existing methods rarely exploit user-side temporal cues and database spatial structure, which could serve as powerful information to improve the stability of VPR.
@@ II-C Fine-grained visual re-localization
For high-accuracy visual re-localization (e.g. decimeter- or centimeter-level), it is generally needed to perform fine-grained association between user perception and the map elements. Many of these methods follow a coarse-to-fine pipeline. In ORB-SLAM3[8] and VINS-Mono[13], re-localization is performed through bag-of-words-based place recognition and point-level matching between feature observations and the point cloud map. Such pipeline is further improved in HLoc[9], where DNN-based models are fully utilized. Besides, monocular depth prediction is exploited in [33] to achieve map-free visual re-localization.
Other approaches try to exploit higher-level information for map matching, including instance-level object modeling[34, 35], bird’s-eye view (BEV) features[36] and so on. The tradeoff of such semantic-based re-localization is the loss of detailed, sometimes subtle but distinguishable, visual cues, which might lead to higher ambiguity.
Moreover, relocalization based on neural/learnable representations (e.g., NeRF, 3DGS) has recently attracted attention. However, existing work usually focuses on small scenes[11, 12] or image pose regression within the same sequence used for model construction[37], lacking practical studies on large-scale map-based geo-localization.
@@ III System Overview
The proposed system is mainly divided into the mapping phase and the localization phase, as shown in Fig. 2.
In the mapping phase, multi-sensor information is used to generate the feature-free map frames with recovered dense depth. Global optimization which incorporates local visual constraints is performed periodically at low frequency to achieve optimal pose estimation. On this basis, the local co-visibility is checked to keep the map sparsity and to achieve incremental mapping. Besides, a DNN-based VPR model is employed to compute the global descriptors of map frames.
In the localization phase, deep VPR and feature matching models are exploited to achieve coarse-to-fine localization. During this process, user-side multi-frame information and the map frame distribution are fully leveraged to improve the recall rate and the accuracy.
@@ IV Mapping Phase
In this section, the implementation details of the mapping phase are presented.
@@ IV-A Multi-Sensor Dense Bundle Adjustment
The foundation of the mapping process is to obtain accurate geometric information, i.e. the absolute poses and dense depths of the images. To achieve this, multi-sensor dense bundle adjustment (MS-DBA) with global optimization is employed.
The concept of dense bundle adjustment (DBA) is proposed in[21]. Specifically, sequential images are fed into a convolutional neural network (CNN) with gated recurrent units (GRUs) to compute dense optical flows among co-visible image pairs. To explain this, consider the rigid flow produced by dense projection of one image pair (source frame $i$ projected to target frame $j$)
@@
where $\Pi_{c}$ is the camera projection model, $\mathbf{T}_{ij}$ is the relative camera pose which lies in the 3D special Euclidean group SE(3), $\circ$” is the transformation operation of SE(3), $\mathbf{u}_{i}$ denotes the $n=(H/8)\times(W/8)$ grid-like 2D points in the image frame which balance scene structure completeness and computational cost, $\bm{\lambda}_{i}$ is the inverse depth map.
The optical flow module takes the rigid flow and the CNN-encoded image feature map as input and outputs the corrected optical flow (in the residual form) together with the weight $\mathbf{w}_{ij}$, leading to the following linearized error equation
$\left(\delta\mathbf{u}_{ij}\right)_{2n\times 1}=\left[\begin{matrix}\mathbf{J}_{i}&\mathbf{J}_{j}&\mathbf{J}_{\bm{\lambda}_{i}}\end{matrix}\right]\left[\begin{matrix}\bm{\xi}_{i}^{\top}&\bm{\xi}_{j}^{\top}&\left({\delta\bm{\lambda}_{i}}\right)_{n\times 1}^{\top}\end{matrix}\right]^{\top}$ (2)
where $\bm{\xi}_{i}$, $\bm{\xi}_{j}$ are Lie algebras of the camera poses (corresponding to $\mathbf{T}_{i}$ and $\mathbf{T}_{j}$), $\mathbf{J}_{i}$, $\mathbf{J}_{j}$, $\mathbf{J}_{\bm{\lambda}_{i}}$ are Jacobians derived from (1).
The residual-form dense optical flows are treated as re-projection errors, which are used for bundle adjustment to simultaneously estimate the poses and dense depths of the images. This process is recurrently performed and is end-to-end trainable, which shows good generalization in real-world scenes with synthetic training[21].
In our early work[40], we demonstrate the possibility of tightly integrating DBA with multi-sensor information. To be specific, the Hessian form of Eq. (2) is derived and stacked for all projections from frame $i$, leading to
$\left[\begin{matrix}\mathbf{v}_{i}\\ \mathbf{z}_{i}\end{matrix}\right]=\left[\begin{matrix}\mathbf{B}_{i}&\mathbf{D}_{i}\\ \mathbf{D}^{\top}_{i}&\mathbf{C}_{i}\end{matrix}\right]\left[\begin{matrix}\bm{\xi}_{i,1,2,\dots,N}\\ {\delta\bm{\lambda}_{i}}\end{matrix}\right]$ (3)
where $\mathbf{B}_{i}$, $\mathbf{C}_{i}$, $\mathbf{D}_{i}$ are block matrices of the Hessian matrix, $\bm{\xi}_{i,1,2,\cdots,N}$ is the stacked Lie algebras of camera poses.
@@
$(\underbrace{\mathbf{B}_{i}-\mathbf{D}_{i}\mathbf{C}_{i}^{-1}\mathbf{D}_{i}^{\top}}_{\mathbf{H}_{c,i}})\bm{\xi}_{i,1,2,\cdots,N}=\underbrace{\mathbf{v}_{i}-\mathbf{D}_{i}\mathbf{C}_{i}^{-1}\mathbf{z}_{i}}_{\mathbf{v}_{c,i}}$ (4)
in which $\mathbf{C}_{i}$ is a diagonal matrix.
@@
The Hessian-form pose constraint Eq. (4), which embeds the DBA information of co-visible images, is fed to a factor graph based on GTSAM[39] for multi-sensor fusion, as shown in Fig. 3. A generic sliding-window graph design is employed, which allows for flexible integration of different sensor information. The states $\mathbf{x}$ in the graph are as follows
$\mathbf{x}=\left\{\mathbf{x}_{k}\mid k\in\mathcal{K}\right\},\ \ \mathbf{x}_{k}=\left(\begin{matrix}\mathbf{T}^{w}_{b_{k}}&\mathbf{v}^{w}_{b_{k}}&\mathbf{b}_{k}\end{matrix}\right)$ (5)
$\mathbf{T}^{w}_{b_{k}}=\left[\begin{matrix}\mathbf{R}^{w}_{b_{k}}&\mathbf{t}^{w}_{b_{k}}\\ 0&1\end{matrix}\right],\ \ \mathbf{b}_{k}=\left(\begin{matrix}\mathbf{b}_{a,k}&\mathbf{b}_{g,k}\end{matrix}\right)$ (6)
where $\mathbf{x}_{k}$ is the state at time $k$, $\mathcal{K}$ represents the set of timestamps within the sliding window, $\mathbf{T}^{w}_{b_{k}}$ is the body (IMU) poses in the world frame, $\mathbf{v}^{w}_{b_{k}}$ is the velocity, $\mathbf{b}_{a,k}$ and $\mathbf{b}_{g,k}$ are accelerometer/gyroscope biases.
The following factors are considered in this work:
@@
$\mathbf{E}_{c}(\mathbf{x})=\frac{1}{2}l_{c}\!\left(\mathbf{x}\right)^{\top}\mathbf{H}_{c}\,l_{c}\!\left(\mathbf{x}\right)-l_{c}\!\left(\mathbf{x}\right)^{\top}\mathbf{v}_{c}$ (7)
where $\mathbf{H}_{c}$, $\mathbf{v}_{c}$ are the Hessian matrix and vector of DBA (stacked from $\mathbf{H}_{c,i}$, $\mathbf{v}_{c,i}$ in Eq. (4)) with all the depth states eliminated, $l_{c}(\cdot)$ is the linear container that transform the current states $\mathbf{x}$ to the Lie algebras of camera poses in Eq. (4) at specific linearization point.
@@
where $\tilde{\mathbf{p}}^{n}_{g}$ is the measured position of GNSS antenna in the navigation frame, $\mathbf{t}^{b}_{g}$ is the lever-arm, $\mathbf{T}^{n}_{w}$ is a fixed world-to-navigation transformation obtained from initial alignment.
Probabilistic marginalization[14] is applicable to the above factor graph based on Schur complement to ensure real-time performance. Thus, the overall cost function of the sliding-window factor graph is defined as follows
$\begin{split}\min\ \ \mathbf{E}_{m}(\mathbf{x})+\mathbf{E}_{c}(\mathbf{x})&\\ +\sum_{k\in\mathcal{K}}\left\|\mathbf{r}_{b}(\mathbf{x}_{k},\mathbf{x}_{k+1})\right\|^{2}_{\bm{\Omega}_{b}}&+\sum_{k\in\mathcal{G}}\rho_{C}\left(\left\|\mathbf{r}_{g}(\mathbf{x}_{k})\right\|^{2}_{\bm{\Omega}_{g}}\right)\end{split}$ (10)
where $\mathbf{E}_{m}(\mathbf{x})$ is the marginalization term in the quadric cost function form, $\mathcal{G}$ is the set of GNSS measurement epochs in the sliding window, $\bm{\Omega}_{b}$ and $\bm{\Omega}_{g}$ are uncertainties of the factors, $\rho_{C}(\cdot)$ is the Cauchy loss function.
@@
$\delta\bm{\lambda}_{i}=\mathbf{C}_{i}^{-1}(\mathbf{z}_{i}-\mathbf{D}_{i}^{\top}\bm{\xi}_{i,1,2,\dots,N})$ (11)
then the estimated poses and depths are fed back into the optical flow module to refine the optical flows $\delta\mathbf{u}_{ij}$ and adjust the weights $\mathbf{w}_{ij}$. The overall process simultaneously adjust the pose estimation, the dense data association and the depths, trying to achieve consistency of the above information.
For mapping tasks, an issue that should be considered is the obtainment of global optimal pose estimation. In realistic scenarios, both the observability degradation of monocular visual-inertial odometry (VIO) and intermittent GNSS outages would lead to the decline of pose accuracy, which can hardly be overcome by a sliding-window estimator. To handle this, a global factor graph is maintained in our system, which collects the marginalized factors from the sliding-window factor graph for further global optimization. The objective function of the global factor graph is written as follows
@@
where $\mathcal{K}_{\text{all}}$ is the set of timestamps in the global graph, $\mathcal{G}_{\text{all}}$ is the set of all GNSS measurement epochs, $\mathbf{E}_{c,k}(\mathbf{x}_{c,k})$ is the marginalized DBA information taking frame $k$ as the source frame (corresponding to Eq. (4)), $\mathbf{x}_{c,k}$ is the related poses. Each $\mathbf{E}_{c,k}$ reflects the local visual constraint around frame $k$.
@@
Figure 4: Illustration of the co-visibility checking. (a) Co-visibility for one image pair; (b) Local co-visibility checking, the dashed-line frame would be added/discarded based on its score comparison with the high-covisbility map frame; (c) An example of the sparsified structure frame map.